OpenStudy (xxstudyingstuffxx):

So, I was taking a math est and it gave me a confusing question, and I spent a half hour trying to solve it, but I still don't understand. I will post the question in the comments.

OpenStudy (518nad):

k

OpenStudy (518nad):

:)

OpenStudy (xxstudyingstuffxx):

Please no direct answers

OpenStudy (518nad):

you got it right already

OpenStudy (518nad):

they should state which shows a proportional relationship as shown in the graph

OpenStudy (518nad):

because any 2 given points are a proportional relationship already

OpenStudy (xxstudyingstuffxx):

I dont understand what you mean.

OpenStudy (xxstudyingstuffxx):

@518nad

OpenStudy (518nad):

okay which part

OpenStudy (xxstudyingstuffxx):

Your last two comments

OpenStudy (518nad):

okay i think i get what ur question is looking for

OpenStudy (518nad):

so they want you to identify points in the form y=mx, this is a strictly proportional relationship y=mx+b is a proportion and shift

OpenStudy (518nad):

the 2 lines given to you on the graph, are in y=mx form because they pass through zero so b=0 y=mx+0 = mx the points you selected from the first choice are points on the graph

OpenStudy (518nad):

(3,6) (5,10)

OpenStudy (xxstudyingstuffxx):

I am sorry. I probably sound dumb. I dont even know what it is asking for (I skipped 7th grade)

OpenStudy (518nad):

its okay, you knw the equation of a line is y=mx+b

OpenStudy (518nad):

m is the slope and b is the y intercept

OpenStudy (xxstudyingstuffxx):

Still dont have a clue what you are saying. :(

OpenStudy (518nad):

the equation of a line is Y=mx+b

OpenStudy (xxstudyingstuffxx):

ok

OpenStudy (518nad):

um okay lets start with y=mx

OpenStudy (518nad):

are you away of how to use the axis

OpenStudy (518nad):

are you aware*

OpenStudy (xxstudyingstuffxx):

yes

OpenStudy (518nad):

okay so a line equation is this expression that tells you how to draw a line on an axis

OpenStudy (518nad):

for example if u have the line equation y=2x

OpenStudy (518nad):

you can plug in a value for x and it will tell you what y is, and this will give u a point (x,y) to graph

OpenStudy (xxstudyingstuffxx):

So if you had the equasion y=2x, and x = 4, then 7 would be 8?

OpenStudy (518nad):

yes y would be 8 then

OpenStudy (518nad):

this is what we call a proportional relationship, when its in the form y=mx, where m is some number

OpenStudy (518nad):

for the set of 2 points you picked in the first choice

OpenStudy (518nad):

both of those 2 points satisfy a proportional relationship

OpenStudy (518nad):

3,6 x,2x 5,10 x,2x so y=2x here

OpenStudy (xxstudyingstuffxx):

Ok

OpenStudy (518nad):

for the other set of points, they do not satisfy, because they have different proportions they are not on the same proportion line

OpenStudy (518nad):

for example 5,10 x,2x y=2x 12,6 x,x/2 y=1/2 * x these are 2 differnt proportion equation one is y=2x, another is y=1/2 * x

OpenStudy (xxstudyingstuffxx):

So what you are saying is that we have to solve 3,6 x, 2x 5,10 x, 2x correct?

OpenStudy (518nad):

ya im showing u the relationship

OpenStudy (xxstudyingstuffxx):

Oh ok

OpenStudy (518nad):

3,6 3,2*3 5,10 5,2*5 the y value is the x being multipled by 2

OpenStudy (xxstudyingstuffxx):

So for 3,6 x, 2x x would be equal to 3, right?

OpenStudy (518nad):

yes

OpenStudy (xxstudyingstuffxx):

So y is 6?

OpenStudy (518nad):

yes

OpenStudy (518nad):

i took 3,6 and said its in the form x,2x since its 2 times the x value

OpenStudy (xxstudyingstuffxx):

So it would be x, 2x = 3,6

OpenStudy (518nad):

yes umm

OpenStudy (518nad):

are you confused with the point notation

OpenStudy (xxstudyingstuffxx):

No, you have helped me understand it alot! Thanks! Sorry for needing such a long explanation.

OpenStudy (518nad):

no problem!

OpenStudy (518nad):

as long as u get it

OpenStudy (518nad):

if both of your points given satisfy the same y=mx equation for some m

OpenStudy (518nad):

then both of the points are on the same proportion line

OpenStudy (518nad):

in this case the first set of points both satisfy the euation y=2x all the other set of points given satisfy different y=mx lines

OpenStudy (xxstudyingstuffxx):

Thanks! I understand it now.

OpenStudy (xxstudyingstuffxx):

Thanks sooooooooo much!

OpenStudy (518nad):

okay good lol :)

OpenStudy (xxstudyingstuffxx):

☺☻☺☻

OpenStudy (518nad):

☺☻